# Characterisation of the big Ramsey degrees of the generic partial order

### v3

@@ 1. Introduction
Poset-diaries of small partial orders can be determined by an exhaustive search tool. [^1] https://github.com/janhubicka/big-ramsey We determined that $|T(\mathbf{C}_{3})|=52$, $|T(\mathbf{C}_{4})|=11000$, $|T(\mathbf{A}_{3})|=84$, and $|T(\mathbf{A}_{4})|=75642$. Overall there are:
@@ 2.2. Trees
Given words $w,w^{\prime}\in\Sigma^{*}$, we write $w\sqsubseteq w^{\prime}$ if $w$ is an initial segment of $w^{\prime}$. With this partial order we obtain a tree $(\Sigma^{*},\sqsubseteq)$ and the notation on words introduced in Section 1 can be viewed as a special case of the notation introduced here.
@@ 3. Tree of 1-types
Poset-diaries, which can be compared to Devlin embedding types (see Chapter 6.3 of [35]) or the diaries introduced in [1], have an intuitive meaning when understood in the context of the tree of 1-types of $\mathbf{P}$. We now introduce this tree and its enrichment to an aged coding tree and discuss how poset-diaries can be obtained as a suitable abstraction of the aged coding tree.
An enumerated structure is simply a structure $\mathbf{A}$ with underlying set $A=|A|$. Fix a countably infinite enumerated structure $\mathbf{A}$. Given vertices $u,v$ and an integer $n$ satisfying $\min(u,v)\geq n\geq 0$, we write $u\sim^{\mathbf{A}}_{n}v$ and say that $u$ and $v$ have the same (quantifier-free) type over $\{0,1,\ldots,n-1\}$, if the structures induced by $\mathbf{A}$ on $\{0,1,\ldots,n-1,u\}$ and $\{0,1,\ldots,n-1,v\}$ are isomorphic via the map which is the identity on $\{0,...,n-1\}$ and sends $u$ to $v$. We write $[u]^{\mathbf{A}}_{n}$ for the $\sim^{\mathbf{A}}_{n}$-equivalence class of vertex $u$.
@@
Let $\mathbf{A}$ be a countably infinite (relational) enumerated structure. Given $n<\omega$, write $\mathbb{T}_{\mathbf{A}}(n)=\omega/\!\sim^{\mathbf{A}}_{n}$. A (quantifier-free) 1-type is any member of the disjoint union $\mathbb{T}_{\mathbf{A}}:=\bigsqcup_{n<\omega}\mathbb{T}_{\mathbf{A}}(n)$. We turn $\mathbb{T}_{\mathbf{A}}$ into a tree as follows. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$ and $y\in\mathbb{T}_{\mathbf{A}}(n)$, we declare that $x\leq^{\mathbb{T}}_{\mathbf{A}}y$ if and only if $m\leq n$ and $x\supseteq y$.
@@
In the case that $\mathbf{A}$ is a structure in a finite binary relational language, we can encode $\mathbb{T}_{\mathbf{A}}$ as a subtree of $k^{<\omega}$ for some $k<\omega$ as follows. Given two enumerated structures $\mathbf{B}$ and $\mathbf{C}$, an ordered embedding of $\mathbf{B}$ into $\mathbf{C}$ is any embedding of $\mathbf{B}$ into $\mathbf{C}$ which is an increasing injection of the underlying sets. Write $\mathrm{OEmb}(\mathbf{B},\mathbf{C})$ for the set of ordered embeddings of $\mathbf{B}$ into $\mathbf{C}$. Fix once and for all an enumeration $\{\mathbf{B}_{i}:i<k\}$ of the set of enumerated structures of size $2$ which admit an enumerated embedding into $\mathbf{A}$. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$, we define $\sigma(x)\in k^{m}$ where given $j<m$, we set $\sigma(x)_{j}=i$ iff for some (equivalently every) $n\in x$ there is $f\in\mathrm{OEmb}(\mathbf{B}_{i},\mathbf{A})$ with $\mathrm{Im}(f)=\{j,n\}$. The map $\sigma\colon\mathbb{T}_{\mathbf{A}}\to k^{<\omega}$ is then an embedding of trees. We write $\mathrm{CT}^{\mathbf{A}}=\sigma[\mathbb{T}_{\mathbf{A}}]$ and call this the coding tree of $\mathbf{A}$. Typically we also endow $\mathrm{CT}^{\mathbf{A}}$ with coding nodes, where for each $n$, the $n^{\mathrm{th}}$ coding node of $\mathrm{CT}^{\mathbf{A}}$ is defined to be $c^{\mathbf{A}}(n):=\sigma([n]^{\mathbf{A}}_{n})$.
@@ 4. Poset-diaries and level structures
Given a poset-diary $S$, one can view $\overline{S}$ as a binary branching tree and each level $\overline{S}_{\ell}$ as a structure $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ where the structure $\mathbf{S}_{\ell+1}$ is constructed from the structure $\mathbf{S}_{\ell}$ as described in the following proposition.
@@
Let $S$ be a poset-diary. Then all words in $\overline{S}$ are mutually compatible, and for each $\ell\leq\sup_{w\in S}|w|$ the structures $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ and $\mathbf{S}_{\ell+1}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ are related as follows:
@@
    $u\preceq v\iff v^{\frown}c\preceq u^{\frown}c^{\prime}$,
@@ 6. Interesting levels and sub-diaries
Fix $\ell\geq 0$, $S=\{u^{0}\mathbin{<_{\mathrm{lex}}}u^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}u^{n-1}\}$, $\Sigma^{*}_{\ell}=\{v^{0}\mathbin{<_{\mathrm{lex}}}v^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}\allowbreak v^{m-1}\}$, and a boring extension $e$ of $S$. For $u\in S$ denote by $i(u)$ the integer $i$ satisfying $u^{i}=u$. For $v\in\Sigma^{*}_{\ell}\setminus S$ we say that letter $c\in\Sigma$ is safe for $v$ if for every $0\leq j<n$ such that $u^{j}$ is compatible with $v$, it holds that ($(\{u^{j},v\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is isomorphic to ($(\{{u^{j}}^{\frown}e_{j},v^{\frown}c\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$.
First we check that for every $v\in\Sigma^{*}\setminus S$, the set of safe letters for $v$ is non-empty. To see this, consider some vertex $v\in\Sigma^{*}\setminus S$ such that $\mathrm{X}$ is not safe for $v$. In this case there are two options:
@@
    There is $u\in S$ compatible with $v$ such that $u\mathbin{<_{\mathrm{lex}}}v$, $u\not\perp v$, and $e_{i(u)}=\mathrm{R}$. In this case we argue that $\mathrm{R}$ is safe for $v$ (in fact, it is the only safe letter for $v$). For this we need to check:
@@
      For every $w\in S$ compatible with $v$ satisfying $v\mathbin{<_{\mathrm{lex}}}w$ and $v\not\perp w$, we have $e_{i(w)}=\mathrm{R}$. This follows from the fact that $v\trianglelefteq w$, and this needs to be preserved by the extension.
@@
      For every $w\in S$ compatible with $v$ satisfying $v\mathbin{<_{\mathrm{lex}}}w$, it holds that $e_{i(w)}\neq\mathrm{L}$. From the condition on $S$, we know that $u$ and $w$ are compatible and that $u\mathbin{<_{\mathrm{lex}}}v\mathbin{<_{\mathrm{lex}}}w$. But then $u^{\frown}\mathrm{R}$ and $w^{\frown}\mathrm{R}$ are incompatible, which contradicts the definition of boring extension.
    3. (c)
      For every $w\in S$ compatible with $v$ satisfying $w\mathbin{<_{\mathrm{lex}}}v$ and $v\not\preceq w$, we have that $e_{i(w)}\neq\mathrm{L}$. Suppose for a contradiction that $e_{i(w)}=\mathrm{L}$. As $w$ and $u$ are compatible and $e$ is boring, we know that $w\prec u$. But, by Proposition 3.6 (P4) and (P6), we know that $w\prec u$, $w\not\prec v$, and $u\perp v$ cannot be all satisfied at the same time, a contradiction.
@@
    There is $u\in S$ compatible with $v$ such that $v\mathbin{<_{\mathrm{lex}}}u$, $v\not\perp u$ and $e_{i(u)}=\mathrm{L}$. In this case we can argue similarly to show that $\mathrm{L}$ is the only safe letter for $v$.
Now we define $e^{\prime}_{j}$ to be $e_{i(v^{j})}$ whenever $v^{j}\in S$; otherwise choose the first letter from $\mathrm{X},\mathrm{L},\mathrm{R}$ that is safe for $v^{j}$.
@@
    $(e^{\prime}_{i},e^{\prime}_{j})\neq(\mathrm{R},\mathrm{L})$. Towards a contradiction, assume that $(e^{\prime}_{i},e^{\prime}_{j})=(\mathrm{R},\mathrm{L})$, and let $u\in S$ be the vertex that made $\mathrm{X}$ unsafe for $v^{i}$ and $u^{\prime}\in S$ be the vertex that made $\mathrm{X}$ unsafe for $v^{j}$. By the same analysis as above we have $u\mathbin{<_{\mathrm{lex}}}v^{i}$, $e_{i(u)}=\mathrm{R}$, $u\not\perp v^{i}$, $v^{j}\mathbin{<_{\mathrm{lex}}}u^{\prime}$, $e_{i(u^{\prime})}=\mathrm{L}$, $v^{j}\not\perp u^{\prime}$. From this however we conclude that $u\mathbin{<_{\mathrm{lex}}}u^{\prime}$, $e_{i(u)}=\mathrm{R}$, $e_{i(u^{\prime})}=\mathrm{L}$ which contradicts the definition of boring extension and our assumption that $u$ and $u^{\prime}$ are compatible.
@@
    $v^{i}\not\perp v^{j}\implies(e^{\prime}_{i},e^{\prime}_{j})\not\in\{(\mathrm{X},\mathrm{L}),(\mathrm{R},\mathrm{X})\}$. Towards a contradiction, assume that $v^{i}\not\perp v^{j}$ and $(e^{\prime}_{i},e^{\prime}_{j})=(\mathrm{X},\mathrm{L})$, and denote by $u\in S$ the vertex that made $\mathrm{X}$ unsafe for $v^{j}$. Clearly $v^{j}\mathbin{<_{\mathrm{lex}}}u$, $v^{j}\not\perp u$ and $e_{i(u)}=L$. It follows that for every $\ell^{\prime}<\ell$ we have $v^{i}_{\ell^{\prime}}\mathbin{\leq_{\mathrm{\Sigma}}}v^{j}_{\ell^{\prime}}\mathbin{\leq_{\mathrm{\Sigma}}}u_{\ell^{\prime}}$. From this we have that $u$ and $v^{i}$ are compatible and $u\not\perp v^{i}$ which makes $\mathrm{X}$ unsafe for $v^{i}$. A contradiction. A very similar argument shows that $(e^{\prime}_{i},e^{\prime}_{j})\neq(\mathrm{R},\mathrm{X})$.
@@
    $v^{i}\not\preceq v^{j}\implies(e^{\prime}_{i},e^{\prime}_{j})\neq(\mathrm{L},\mathrm{R})$. Towards a contradiction, assume that $v^{i}\not\preceq v^{j}$ and $(e^{\prime}_{i},e^{\prime}_{j})=(\mathrm{L},\mathrm{R})$, and let $u\in S$ be the vertex that made $\mathrm{X}$ unsafe for $v^{i}$ and $u^{\prime}\in S$ be the vertex that made $\mathrm{X}$ unsafe for $v^{j}$. We have $v^{i}\mathbin{<_{\mathrm{lex}}}u\mathbin{<_{\mathrm{lex}}}u^{\prime}\mathbin{<_{\mathrm{lex}}}v^{j}$ and $e_{i(u)}=\mathrm{L}$, $e_{i(u^{\prime})}=\mathrm{R}$. It follows that $v^{i}\not\perp u$, $u\preceq u^{\prime}$, $u^{\prime}\not\perp v^{j}$. Now for every $\ell^{\prime}<\ell$ we also have $v^{i}_{\ell^{\prime}}\mathbin{<_{\mathrm{\Sigma}}}u_{\ell^{\prime}}\mathbin{<_{\mathrm{\Sigma}}}u^{\prime}_{\ell^{\prime}}\mathbin{<_{\mathrm{\Sigma}}}v^{j}_{\ell^{\prime}}$, which implies that $v^{i}\preceq v^{j}$. A contradiction. ∎

### v4

@@ 1. Introduction
Poset-diaries of small partial orders can be determined by an exhaustive search tool. [^1] https://github.com/janhubicka/big-ramsey We determined that $|T(\mathbf{C}_{3})|=52$, $|T(\mathbf{C}_{4})|=11000$, $|T(\mathbf{A}_{3})|=84$, and $|T(\mathbf{A}_{4})|=75672$. Overall there are:
@@ 2.2. Trees
Given words $w,w^{\prime}\in\Sigma^{*}$, we write $w\sqsubseteq w^{\prime}$ if $w$ is an initial segment of $w^{\prime}$. With this partial order we obtain the tree $(\Sigma^{*},\sqsubseteq)$ and the notation on words introduced in Section 1 can be viewed as a special case of the notation introduced here.
@@ 3. Tree of 1-types
Poset-diaries, which can be compared to Devlin embedding types (see Chapter 6.3 of [41]) or diaries introduced in [3], have an intuitive meaning when understood in the context of the tree of 1-types of $\mathbf{P}$. We now introduce this tree and its enrichment to an aged coding tree and discuss how poset-diaries can be obtained as a suitable abstraction of the aged coding tree.
An enumerated structure is simply a structure $\mathbf{A}$ with underlying set $A=|A|$. Fix a countably infinite enumerated structure $\mathbf{A}$. Given vertices $u,v$ and an integer $n$ satisfying $\min(u,v)\geq n\geq 0$, we write $u\sim^{\mathbf{A}}_{n}v$ and say that $u$ and $v$ have the same (quantifier-free) type over $\{0,1,\ldots,n-1\}$, if the structures induced by $\mathbf{A}$ on $\{0,1,\ldots,n-1,u\}$ and $\{0,1,\ldots,n-1,v\}$ are isomorphic via the map which is the identity on $\{0,...,n-1\}$ and sends $u$ to $v$. We write $[u]^{\mathbf{A}}_{n}$ for the $\sim^{\mathbf{A}}_{n}$-equivalence class of the vertex $u$.
@@
Let $\mathbf{A}$ be a countably infinite (relational) enumerated structure. Given $n<\omega$, write $\mathbb{T}_{\mathbf{A}}(n)=\omega/\!\sim^{\mathbf{A}}_{n}$. A (quantifier-free) 1-type is any member of the disjoint union $\mathbb{T}_{\mathbf{A}}:=\bigsqcup_{n<\omega}\mathbb{T}_{\mathbf{A}}(n)$. We turn $\mathbb{T}_{\mathbf{A}}$ into a tree as follows. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$ and $y\in\mathbb{T}_{\mathbf{A}}(n)$, we declare that $x\leq^{\mathbb{T}}_{\mathbf{A}}y$ if and only if $m\leq n$ and $x\supseteq y$. We will denote by $[n]^{\mathbf{A}}_{n}$ the equivalence class of $n$ under $\sim^{\mathbf{A}}_{n}$.
@@
In the case that $\mathbf{A}$ is a structure in a finite binary relational language, we can encode $\mathbb{T}_{\mathbf{A}}$ as a subtree of $k^{<\omega}$ for some $k<\omega$ as follows. Given two enumerated structures $\mathbf{B}$ and $\mathbf{C}$, an ordered embedding of $\mathbf{B}$ into $\mathbf{C}$ is any embedding of $\mathbf{B}$ into $\mathbf{C}$ which is an increasing injection of the underlying sets. Write $\mathrm{OEmb}(\mathbf{B},\mathbf{C})$ for the set of ordered embeddings of $\mathbf{B}$ into $\mathbf{C}$. Fix once and for all an enumeration $\{\mathbf{B}_{i}:i<k\}$ of the set of enumerated structures of size $2$ which admit an enumerated embedding into $\mathbf{A}$. Given $x\in\mathbb{T}_{\mathbf{A}}(m)$, we define $\sigma(x)\in k^{m}$ (a word in alphabet $\{0,1,\ldots,k-1\}$ of length $m$) where given $j<m$, we set $\sigma(x)_{j}=i$ (recall that we use subscripts to index letters in a word) if and only if for some (equivalently every) $n\in x$ there is $f\in\mathrm{OEmb}(\mathbf{B}_{i},\mathbf{A})$ with $\mathrm{Im}(f)=\{j,n\}$. The map $\sigma\colon\mathbb{T}_{\mathbf{A}}\to k^{<\omega}$ is then an embedding of trees. We write $\mathrm{CT}^{\mathbf{A}}=\sigma[\mathbb{T}_{\mathbf{A}}]$ and call this the coding tree of $\mathbf{A}$. Typically we also endow $\mathrm{CT}^{\mathbf{A}}$ with coding nodes, where for each $n$, the $n^{\mathrm{th}}$ coding node of $\mathrm{CT}^{\mathbf{A}}$ is defined to be $c^{\mathbf{A}}(n):=\sigma([n]^{\mathbf{A}}_{n})$.
@@
Intuitively, the words $u$ and $v$ are compatible if and only if they can appear in a coding tree of some enumeration of $\mathbf{P}$. If there exists $\ell<\min(|u|,|v|)$ such that $(u_{\ell},v_{\ell})=(\mathrm{R},\mathrm{L})$, then in this enumeration we must have $|u|\prec\ell\prec|v|$. Failures of condition (2) of Definition 3.3 contradict transitivity of $\mathbf{P}$.
@@ 4. Poset-diaries and level structures
Given a poset-diary $S$ (Definition 1.5), one can view $\overline{S}$ as a binary branching tree and each level $\overline{S}_{\ell}$ as a structure $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ where the structure $\mathbf{S}_{\ell+1}$ is constructed from the structure $\mathbf{S}_{\ell}$ as described in the following proposition.
@@
Let $S$ be a poset-diary. Then all words in $\overline{S}$ are mutually compatible, and for each $\ell\leq\sup_{w\in S}|w|$ the structures $\mathbf{S}_{\ell}=(\overline{S}_{\ell},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ and $\mathbf{S}_{\ell+1}=(\overline{S}_{\ell+1},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ are related as follows:
@@
    $u\preceq v\iff u^{\frown}c\preceq v^{\frown}c^{\prime}$,
@@ 6. Interesting levels and sub-diaries
Fix $\ell\geq 0$, $S=\{u^{0}\mathbin{<_{\mathrm{lex}}}u^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}u^{n-1}\}$, $\Sigma^{*}_{\ell}=\{v^{0}\mathbin{<_{\mathrm{lex}}}v^{1}\mathbin{<_{\mathrm{lex}}}\cdots\mathbin{<_{\mathrm{lex}}}\allowbreak v^{m-1}\}$, and a boring extension $e$ of $S$. For $u\in S$ denote by $i(u)$ the integer $i$ satisfying $u^{i}=u$. For $v\in\Sigma^{*}_{\ell}\setminus S$ we say that letter $c\in\Sigma$ is safe for $v$ if for every $0\leq j<n$ such that $u^{j}$ is compatible with $v$, it holds that ${u^{j}}^{\frown}e_{j}$ is compatible with $v^{\frown}c$ and ($(\{u^{j},v\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$ is isomorphic to ($(\{{u^{j}}^{\frown}e_{j},v^{\frown}c\},\allowbreak\mathbin{\leq_{\mathrm{lex}}}\nobreak,\allowbreak\preceq\nobreak,\allowbreak\trianglelefteq)$.
First we check that for every $v\in\Sigma^{*}\setminus S$, the set of safe letters for $v$ is non-empty. To see this, consider some vertex $v\in\Sigma^{*}\setminus S$ such that $\mathrm{X}$ is not safe for $v$. Let $u\in S$ witness that $\mathrm{X}$ is not safe for $v$. There are two cases depending on if $u\mathbin{<_{\mathrm{lex}}}v$ or $v\mathbin{<_{\mathrm{lex}}}u$.
@@
    If $u\mathbin{<_{\mathrm{lex}}}v$, then we must have that $u\trianglelefteq v$, but $u^{\frown}e_{i(u)}\not\trianglelefteq v^{\frown}\mathrm{X}$, implying that $e_{i(u)}=\mathrm{R}$. We will show that $\mathrm{R}$ is safe for $v$. Towards a contradiction, suppose that $w\in S$ witnessed that $\mathrm{R}$ was not safe for $v$.
@@
      If $v\mathbin{<_{\mathrm{lex}}}w$, there are two possibilities.
      • $v\trianglelefteq w$, but $v^{\frown}\mathrm{R}\not\trianglelefteq w^{\frown}e_{i(w)}$. This implies that $e_{i(w)}\neq\mathrm{R}$. However, since $u\trianglelefteq w$, we must have $u^{\frown}\mathrm{R}\trianglelefteq w^{\frown}e_{i(w)}$, which implies $e_{i(w)}=\mathrm{R}$, a contradiction.
      • $\{v^{\frown}\mathrm{R},w^{\frown}e_{i(w)}\}$ are incompatible. If $v\trianglelefteq w$, this reduces to the case above. If $v\not\trianglelefteq w$, then we must have $e_{i(w)}=\mathrm{L}$. But as $\{u^{\frown}\mathrm{R},w^{\frown}\mathrm{L}\}$ are incompatible, this is a contradiction.
@@
      If $w\mathbin{<_{\mathrm{lex}}}v$, the only possibility is that $w\not\prec v$, but $w^{\frown}e_{i(w)}\prec v^{\frown}\mathrm{R}$. This implies that $e_{i(w)}=\mathrm{L}$ and that $w\trianglelefteq v$. We cannot have $u\mathbin{<_{\mathrm{lex}}}w$ since $\{u^{\frown}R,w^{\frown}L\}$ would not be compatible, and we cannot have $u=w$ since $\mathrm{R}\neq\mathrm{L}$. So $w\mathbin{<_{\mathrm{lex}}}u$; since $e_{i(w)}=\mathrm{L}$ and $e_{i(u)}=\mathrm{R}$, we must have $w\trianglelefteq u$ (otherwise $\{w^{\frown}\mathrm{L},u^{\frown}R\}$ would not be compatible) and $w\prec u$ (since $w^{\frown}\mathrm{L}\prec u^{\frown}\mathrm{R}$), contradicting that $u\trianglelefteq v$ and $w\not\prec v$.
@@
    If $v\mathbin{<_{\mathrm{lex}}}u$, the argument is very similar to the above; one shows that $\mathrm{L}$ is safe for $v$.
Now we define $e^{\prime}_{j}$ to be $e_{i(v^{j})}$ whenever $v^{j}\in S$; otherwise choose the first letter from $\mathrm{X},\mathrm{L},\mathrm{R}$ that is safe for $v^{j}$ (equivalently from $\mathrm{X},\mathrm{R},\mathrm{L}$, since if $\mathrm{X}$ is not safe, the argument above shows that exactly one of $\mathrm{L}$ or $\mathrm{R}$ is safe).
@@
    $(e^{\prime}_{i},e^{\prime}_{j})\neq(\mathrm{R},\mathrm{L})$. Towards a contradiction, assume that $(e^{\prime}_{i},e^{\prime}_{j})=(\mathrm{R},\mathrm{L})$, and let $u\in S$ be some vertex that made $\mathrm{X}$ unsafe for $v^{i}$ and $u^{\prime}\in S$ be some vertex that made $\mathrm{X}$ unsafe for $v^{j}$. By the same analysis as above we have $u\mathbin{<_{\mathrm{lex}}}v^{i}$, $e_{i(u)}=\mathrm{R}$, $u\not\perp v^{i}$, $v^{j}\mathbin{<_{\mathrm{lex}}}u^{\prime}$, $e_{i(u^{\prime})}=\mathrm{L}$, $v^{j}\not\perp u^{\prime}$. From this however we conclude that $u\mathbin{<_{\mathrm{lex}}}u^{\prime}$, $e_{i(u)}=\mathrm{R}$, $e_{i(u^{\prime})}=\mathrm{L}$ which contradicts the definition of boring extension and our assumption that $u$ and $u^{\prime}$ are compatible.
@@
    $v^{i}\not\perp v^{j}\implies(e^{\prime}_{i},e^{\prime}_{j})\not\in\{(\mathrm{X},\mathrm{L}),(\mathrm{R},\mathrm{X})\}$. Towards a contradiction, assume that $v^{i}\not\perp v^{j}$ and $(e^{\prime}_{i},e^{\prime}_{j})=(\mathrm{X},\mathrm{L})$, and denote by $u\in S$ some vertex that made $\mathrm{X}$ unsafe for $v^{j}$. Clearly $v^{j}\mathbin{<_{\mathrm{lex}}}u$, $v^{j}\not\perp u$ and $e_{i(u)}=L$. It follows that for every $\ell^{\prime}<\ell$ we have $v^{i}_{\ell^{\prime}}\mathbin{\leq_{\mathrm{\Sigma}}}v^{j}_{\ell^{\prime}}\mathbin{\leq_{\mathrm{\Sigma}}}u_{\ell^{\prime}}$. From this we have that $u$ and $v^{i}$ are compatible and $u\not\perp v^{i}$ which makes $\mathrm{X}$ unsafe for $v^{i}$. A contradiction. A very similar argument shows that $(e^{\prime}_{i},e^{\prime}_{j})\neq(\mathrm{R},\mathrm{X})$.
@@
    $v^{i}\not\preceq v^{j}\implies(e^{\prime}_{i},e^{\prime}_{j})\neq(\mathrm{L},\mathrm{R})$. Towards a contradiction, assume that $v^{i}\not\preceq v^{j}$ and $(e^{\prime}_{i},e^{\prime}_{j})=(\mathrm{L},\mathrm{R})$, and let $u\in S$ be some vertex that made $\mathrm{X}$ unsafe for $v^{i}$ and $u^{\prime}\in S$ be the vertex that made $\mathrm{X}$ unsafe for $v^{j}$. We have $v^{i}\mathbin{<_{\mathrm{lex}}}u\mathbin{<_{\mathrm{lex}}}u^{\prime}\mathbin{<_{\mathrm{lex}}}v^{j}$ and $e_{i(u)}=\mathrm{L}$, $e_{i(u^{\prime})}=\mathrm{R}$. It follows that $v^{i}\not\perp u$, $u\preceq u^{\prime}$, $u^{\prime}\not\perp v^{j}$. Now for every $\ell^{\prime}<\ell$ we also have $v^{i}_{\ell^{\prime}}\mathbin{<_{\mathrm{\Sigma}}}u_{\ell^{\prime}}\mathbin{<_{\mathrm{\Sigma}}}u^{\prime}_{\ell^{\prime}}\mathbin{<_{\mathrm{\Sigma}}}v^{j}_{\ell^{\prime}}$, which implies that $v^{i}\preceq v^{j}$. A contradiction. ∎